# Q-Adapter: Customizing Pre-trained LLMs to New Preferences with Forgetting Mitigation

### v3

@@ B.1 Base Model
B.2 Q-Adapter
As introduced in the main body of our paper, Q-Adapter takes advantage of the base model and the downstream preference with a specially designed adapter module. We illustrate the inference-time computational flow of Q-Adapter in Figure 5, where Q-Adapter takes two portions of model outputs to compose the final policy. In Table 3, we list the basic hyper-parameters of Q-Adapter. We format all the training data to the chat format [^6] https://huggingface.co/docs/transformers/main/en/chat_templating to better utilize special tokens in the Llama-3.1 tokenizer (Dubey et al., 2024). We adopt the Hugging Face trainer to control the training procedure with a rewritten loss function, where other hyper-parameters follow the default config of the trainer. We train Q-Adapter in one single machine with 4x NVIDIA GeForce RTX 4090 GPUs. The typical training time cost with 3 epochs is about 10 hours. In inference time, we can derive the final policy $\tilde{\pi}^{*}$ as defined in Equation 4 with a few lines of code like below:
@@ B.2 Q-Adapter
A shortcoming may occur when we apply the above forward function that we cannot directly utilize the past key and values of the adapter model to speed up the inference time, since Q-Adapter requires both the base model output $C_{1}$ and the adapter output $C_{2}$. Nevertheless, we can develop a trivial way to record both outputs $(C_{1},C_{2})$ jointly when caching. In this way we can accelerate Q-Adapter inference by separately feeding the cache into two forward functions, resulting in moderate inference cost like other methods.
B.3 Baselines
@@
We use the TRL (von Werra et al., 2020) implementation of the DPO trainer to train our DPO model with LoRA. According to the documentation of TRL, we optimize the LoRA parameters while keeping the base model as the reference policy. The hyper-parameters of DPO are same as the default of the provided DPO config in TRL.
@@
Appendix C Details of Evaluations
C.1 LLM Benchmarks
@@
C.2 LLM as the Judge
We introduce LLM judgment from alpaca-eval for data where objective metrics are not available. alpaca-eval (Dubois et al., 2024) is an automatic evaluation tool designed to assess the performance of language models on instruction-following tasks. It provides a fast, cost-effective alternative to human evaluation by using high-agreement, GPT-based annotators to compare model outputs with reference responses. It is validated on a large dataset with over 20,000 human-annotated preferences, ensuring strong correlation with human judgment. Since alpaca-eval is not intended for high-stakes scenarios requiring detailed human oversight, we carefully rephrase the prompt of its annotator to better fit our evaluation requirements. For example, a evaluation prompt for the domain of academy in DSP is shown below, where we inject prompts to ask the model to concern more about the academy-related style.
@@ C.2 LLM as the Judge
Appendix D Sensitivity Analysis on $\beta$
In Q-Adapter, we add an additional regularization loss term in an L2 form to constraint the estimate reward range with a hyper-parameter $\beta$. We find that adding such a regularization term can stabilize the training process and thus result in better performance. In Figure 6, we show the MMLU scores of Q-Adapter with different choices of $\beta$. The results demonstrate that introducing the regularization leads to better performance in MMLU. On the other hand, the value of $\beta$ is not very sensitive from $0.01$ to $1.0$. In our main experiments, we train Q-Adapter with $\beta=0.1$.

### v4

@@ C.2 Q-Adapter
As introduced in the main body of our paper, Q-Adapter takes advantage of the base model and the downstream preference with a specially designed adapter module. We illustrate the inference-time computational flow of Q-Adapter in Figure 6, where Q-Adapter takes two portions of model outputs to compose the final policy. In Table 3, we list the basic hyper-parameters of Q-Adapter. We format all the training data to the chat format [^5] https://huggingface.co/docs/transformers/main/en/chat_templating to better utilize special tokens in the Llama-3.1 tokenizer (Team, 2024b). We adopt the Hugging Face trainer to control the training procedure with a rewritten loss function, where other hyper-parameters follow the default config of the trainer. We train Q-Adapter in one single machine with 4x NVIDIA GeForce RTX 4090 GPUs. The typical training time cost with 3 epochs is about 10 hours. In inference time, we can derive the final policy $\tilde{\pi}^{*}$ as defined in Equation 6 with a few lines of code like below:
@@
A shortcoming may occur when we apply the above forward function that we cannot directly utilize the past keys and values of the adapter model to speed up the inference time, since Q-Adapter requires both the base model output and the adapter output. Nevertheless, we can develop a trivial way to record past keys and values from both the base model and the adapter jointly when caching. In this way we can accelerate Q-Adapter inference by separately feeding the cache into two forward functions, resulting in moderate inference cost like other methods. After applying the caching techniques, we can save 25% of the inference time compared to simply forwarding twice, using a single GPU for inference via HuggingFace generation API.
C.3 Baselines
@@
We use the TRL (von Werra et al., 2020) implementation of the DPO trainer to train our DPO model with LoRA. According to the documentation of TRL, we optimize the LoRA parameters while keeping the base model as the reference policy. The hyper-parameters of DPO are same as the default of the provided DPO configuration in TRL.
@@
Appendix D Details of Evaluations
D.1 LLM Benchmarks
@@
Table 4: Length-controlled win rates of Q-Adapter, Replay, PR (DPO), and PR (PPO) in data from the academy and entertainment domains from DSP against the SFT baseline.
Dataset Q-Adapter Replay DPO PPO
GPT-4o
DSP-Academy 55.59 54.57 53.93 53.38
DSP-Entertainment 63.77 46.66 46.27 56.70
Deepseed-chat
DSP-Academy 54.69 54.25 54.34 53.41
DSP-Entertainment 64.13 52.17 49.83 54.11
Claude-3.5-Sonnet
DSP-Academy 53.86 52.07 52.26 53.59
DSP-Entertainment 63.66 51.59 48.63 53.84
D.2 LLM as the Judge
We introduce LLM judgment from AlpacaEval for data where objective metrics are not available. A (Dubois et al., 2024) is an automatic evaluation tool designed to assess the performance of language models on instruction-following tasks. It provides a fast, cost-effective alternative to human evaluation by using high-agreement, GPT-based annotators to compare model outputs with reference responses. It is validated on a large dataset with over 20,000 human-annotated preferences, ensuring strong correlation with human judgment. Since AlpacaEval is not intended for high-stakes scenarios requiring detailed human oversight, we carefully rephrase the prompt of its annotator to better fit our evaluation requirements. For example, a evaluation prompt for the domain of academy in DSP is shown below, where we inject prompts to ask the model to concern more about the academy-related style.
@@ D.2 LLM as the Judge
Our LLM evaluation is auto-evaluated by AlpacaEval (Dubois et al., 2024), where an evaluator of GPT-4 is the de facto configuration and widely adopted by many works. The used length-controlled win rate has 0.98 correlation with humans as reported. To test whether the GPT-4o model may induce potential model bias, we further introduce two powerful models, Claude-3.5-Sonnet (Anthropic, 2024) and Deepseek-chat (Liu et al., 2024a). We use their provided APIs to call the models, where the evaluation process is also via the AlpacaEval framework. We show the length-controlled win rates in Table 4. We re-evaluate the generation texts in two categories of the DSP data along with new baselines including KTO, IPO, and NEFTune. We find that the results among different LLM evaluators are consistent, indicating that Q-Adapter performs well in downstream data.